A parallel-plate capacitor has an area A = 3.00 cm2
and a plate separation d =1 mm. (a) Find
its capacitance. (b) How much charge is on the positive plate if the capacitor is connected to a
6.00-V battery? Calculate (c) the charge density on the positive plate, assuming the density is
uniform, and (d) the magnitude of the electric field between the plates
1
Expert's answer
2020-07-10T10:33:46-0400

As per the given question,

Plate area of the capacitor "(A)=3.0 cm^2=3.0\times 10^{-4}m^2"

Distance between the plates"(d)= 1mm=1\times 10^{-3}m"

Permittivity of free space "(\epsilon_o) =8.85\times 10^{-12}F/m"

i)

We know that the capacitance of the capacitor "=\frac{\epsilon_o A}{d}"

Now, substituting the values,

"C=\frac{8.85\times 10^{-12}\times 3\times 10^{-4}}{1\times 10^{-3}}F"

"\Rightarrow C=26.46\times 10^{-13}F"

"\Rightarrow C= 2.646\times 10^{-12}F"


ii) V=6.0 Volt

Hence charge stored in the capacitor "(q)= cV =2.646\times 10^{-12}\times6 =15.876\times 10^{-12}C"


iii) Charge density on the positive plate of the capacitor "=q\/A"

Substituting the values,


"=\frac{15.876\times 10^{-12}}{3.0\times 10^{-4}}C/m^2"


"=5.292\times 10^{-8}C/m^2"


iv) Magnitude of the electric field between the plates "E=\frac{-\Delta V}{d}"

"=\frac{6}{3\times 10^{-4}}N/C =2\times 10^{4}N/C"